in an auditorium 300 tickets were sold. the total sale pf tickets was rs 1250. if the tickets were of two denominations of rs2.50 And rs5.how much tickets of each denomination were sold?
Answer by JulietG(1812)

If all 300 tickets sold were rs5, the total sale would be 300*5, or 1500.
The actual sale was 1250. 1500-1250 = 250 less
250/rs2.50 = 100
100 tickets were sold at rs2.50 for a total of 250
200 tickets were sold at rs5 for a total of 1000.
250 + 100 = 1250